Afternoon Checkout

23:37 Powering off the adsec and executing and adsc_stop, start

Power On
Connect to AOS
Load Program
SetFlatAO - 638 modes

23:43 Power up the wfs

executing indivisual stepps to see the indivisdual steps and checking communications
w_stop all
w_start_all (everything came up first try and no communication issues to the ad sec)
Operate - chevron on OCAM
Cylcing BCU2K and rebinning. Ok now

Authorizing from IIF to Lucifer. Set focal station on Adsec. Sending preset through AOS. Communication through AOS and WFS. Closing loop with zero gain.

WFS and adsec communicating. WFSCounter updating. LUCI AO checkout good.

1:24 LBTI Checkout
Cycled power on 5, 8, 15, 16
processes already off,
w_start_all All processes up and connected to Msgd

Reauthorization to LBTI
Preset AO - the preset AO updated the focal station, preset AO was a success.
Communication between AOS, Adsec, and WFS verified.

Opened with "Restore last shape", we have a flat now.


Xianyu has a backup version of the adsec software prepared should we want ot go that way.

AO onky 20210530

2:38 Steve is opening and executing a pointing check. Dave is operating LUCI for the speckle test. Steve has prepared a target for OSCO BD+08 2883
2:57 Pointing and collimation complete. Authorizing LUCI mono in preparation for OSCO

3:01 OSCO script sent. Seeing 1.12" on the guider, estimated WFS mag 8.59. Bin 1 1700 Hz
Optimized 3.42 0.54 0.37, apply optical gain failed to converge first try. Really ugly modal gain plot. Risiduals sawtoothing into open loop.

Optimizing again...looks comparable. Dropping H02 down produces under driving so we will live with the orifginal. NCPA on. The system did not release to luci. Aborted luci script, took the first exposure manually and skipped to second item.

TN#031717, 031731, 031743 with NCPA on first position. Seeing varied from 1.2-1.53"
TN#032057, 032111, 032123 with NCPA on second position. Seeing ~1.3-1.5"

TN#032505,1 032518, 032530 with NCPA off first position. DIMM 1.2-1.3"
TN#032814, 032826, 032839 with NCPA off second position. Seeing ~1.05-1.1.5"

3:37 Preset to 11.98 mag targetff or full AO speckle test. Estimated mag 13.33 on wfs from check flux read through AOS. Bin2 912 Hz
Seeing 1.12" on the DIMM
Optimized. Bin2 went to crazy ranges, tt up to 70 during optimization
1.35" on DIMM, 1" on the right guider.
TN#034753, TN#034940,

Opened loop after coming back from sky offset. Just off of hotspot. 26.05. 3.81 2.16,

4:04 stopAO in preparation for ESM script to the same target. Seeing 1.08" on the DIMM. Optimize gain for ESM bin2 really takes too long. 8.8,12.55, 0.08. executes HO1 optimization in both rounds... TN 041520(offsetting), TN041542, TN041720

4:40 Thick clouds, lost the guide star completely.

4:53 Moving to Hoag's Object. TTM on the left, active on the right. Heavy cloud cover.

LBTI checkout started

5:41 Pointing to alpha boo. Soft NCPA on. Closed ok

TN055640, 055903, 060169 might have caught a not-a-safeskip where teh adsec does not increment a safeskip but ocam looks like its safeskipping.

Running with scripts similar to yesterday. Seeing 1.6"

6:08 1 CA error so far on the adsec

6:09. Seinding to a 6th mag HD131951. Soft NCPA

6:19 Moving to star of comparable brightness to yesterdays targets, ~8.9 on the WFS HD131961 pupils oscillating, gopt sucky but working, TN 061917, 062133.

6:24:56 A batch of CA errors in the BCU log on the adsec. OCAM not live. Adsec went into failure. Lmircam was not taking data at time of failure, we were stationary on target, closed loop when the failure occurred.

6:33 core dump of msgd

Adsc stop start, setflat Ao flat set with 577 modes at 57 degrees in elevation.

After this failure, based on discussions thisafternoon, Xiany and Guido will spend time reviewing logs and Ao operations will be ceased intil further notice. Engineering will take over.



-- JenniferPower - 30 May 2021
Topic revision: r8 - 31 May 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback